OpenStudy (anonymous):

In the xy- plane, line q and r are parallel. Line q passes through the point (3, 7), and line r passes through the point (3, 2). If the y-intercept for line r is -2, what is the equation of a line perpendicular to line q, which intersects line q on the y-axis? (a) y= 4/3x + 3 (b) y= -3/4x + 3 (c) y= -4/3x + 11 (d) y= 4/3x + 11 (e) -3/4x + 11

OpenStudy (helder_edwin):

since q and r are parallel lines they have the same slope. so \[ \large \text{q:}\qquad y-7=m(x-3) \] \[ \large \text{r:}\qquad y-2=m(x-3) \]

OpenStudy (helder_edwin):

ok?

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

for line r u r told that the y-intercept is -2. this means that when x=0 then y=-2

OpenStudy (anonymous):

alright

OpenStudy (helder_edwin):

in the equation for r replace x=0 and y=-2 and find the value of m. try it!

OpenStudy (anonymous):

wait what's the equation?

OpenStudy (helder_edwin):

i posted it before

OpenStudy (anonymous):

oh sorry yeah

OpenStudy (helder_edwin):

it is my first post

OpenStudy (anonymous):

-4 = m(0-3)

OpenStudy (helder_edwin):

yes. find m

OpenStudy (anonymous):

4/3 ?

OpenStudy (helder_edwin):

great

OpenStudy (helder_edwin):

this means that the equation for line q is \[ \large \text{q:}\quad y-7=\frac{4}{3}(x-3) \]

OpenStudy (helder_edwin):

ok?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

it can be rewrittren into \[ \large y-7=\frac{4}{3}x-4 \] or \[ \large y=\frac{4}{3}x+3 \] (slope-intercept form)

OpenStudy (helder_edwin):

got it?

OpenStudy (anonymous):

yeah thanks

OpenStudy (helder_edwin):

now finally. u r asked for line perpendicular to this one that passes through the y-intercept of q (its intersection with the y-axis)

OpenStudy (helder_edwin):

the y-intercept of q is (0,4)

OpenStudy (helder_edwin):

q has slope m=4/3 the new line has slope m1 such that \[ \large m_1\times m=-1 \] what would be the value of m1?

OpenStudy (anonymous):

is the new slope -3/4

OpenStudy (helder_edwin):

yes so the line requested would be \[ \large y-4=-\frac{3}{4}(x-0) \]

OpenStudy (helder_edwin):

sorry. made a mistake

OpenStudy (helder_edwin):

the y-intercept is (0,3)

OpenStudy (helder_edwin):

then the equation is \[ \large y-3=-\frac{3}{4}(x-0) \]

OpenStudy (helder_edwin):

got it?

OpenStudy (helder_edwin):

hope u got it. gotta go. bye

OpenStudy (anonymous):

thanks got it (y)